OpenStudy (anonymous):

What is the value of x, given that OP ll NQ? a. x=10 b. x=20 c. x=13 d. x=25.5

OpenStudy (anonymous):

Part of your question is definetely missing.

OpenStudy (anonymous):



OpenStudy (anonymous):

Since OP is parallel to NQ, you've got similar triangles. So: \[\frac{ 20 }{ 20+x }=\frac{ 18 }{ 9+18 }\]

OpenStudy (anonymous):

ok next step please im lost

OpenStudy (anonymous):

NM ii only have seven minutes left

OpenStudy (anonymous):

-_- x=10